# Gubkin str. 8, 119991 Moscow, Russia

### v1

@@ Abstract
We study a Euclidean quantum field theory of two interacting real scalar fields in $D=4-\epsilon$ dimensions with the most general two–derivative but Lorentz–violating quadratic structure, allowing anisotropic spatial gradients and field–mixing time and cross–gradient terms, together with general cubic and quartic interactions. Although Lorentz violation is introduced only through the kinetic sector, the renormalization structure is nontrivial: interactions generate additional couplings, including masses and linear terms, so a consistent renormalization–group analysis cannot be formulated in terms of modified kinetic terms alone and requires an RG–complete operator basis compatible with the reduced symmetry.
We perform a systematic one–loop renormalization and derive the complete set of beta functions for cubic and quartic couplings, masses, and linear terms. We identify and analyze the resulting fixed points and fixed manifolds, showing in particular how anisotropy restricts their existence and stability. In particular we obtain how anisotropy restricts existence of canonical Wilson-Fisher fixed point. Also we provide a transparent physical interpretation of the anisotropy–dependent coefficients appearing in the beta functions and clarify how kinetic mixing reshapes the interaction flow through the available ultraviolet phase space.
@@ 1 Introduction
Lorentz invariance is one of the organizing principles of high–energy physics: it underlies the locality and analyticity properties of relativistic QFT, it severely constrains the spectrum of operators, and it provides the kinematical backbone for essentially all perturbative calculations. At the same time, there are many reasons—both conceptual and practical—to study controlled departures from exact Lorentz symmetry. On the one hand, Lorentz invariance may be an emergent symmetry of an interacting many–body system at long distances: in condensed–matter realizations, different fields typically propagate with different “speeds” and the UV theory is anisotropic, while approximate relativistic invariance can appear only in the deep IR Chadha:1982ex ; Volovik:2003fe ; Wen:2004ym . On the other hand, in particle physics and quantum gravity it is natural to ask how robust the Lorentz–invariant fixed points are under Lorentz–violating perturbations, and whether small violations can remain technically natural under renormalization Myers:2003fd ; Collins:2004bp ; GrootNibbelink:2004za (as explored, for instance, in EFT frameworks such as the Standard–Model Extension and its gravitational generalizations Kostelecky:1989jp ; Colladay:1996iz ; Colladay:1998fq ; Kostelecky:2003fs ; Kostelecky:2008ts and in the broader phenomenological literature Mattingly:2005re ; Jacobson:2005bg ; Liberati:2013xla ).
@@
There are (at least) two complementary ways in which Lorentz invariance can be violated in a continuum field theory. The first is to change the scaling structure by introducing higher spatial derivatives (Lifshitz–type theories with dynamical exponent $z\neq 1$, including the Hořava–Lifshitz construction in gravity) Lifshitz:1941 ; Horava:2009uw ; Sotiriou:2009gy ; Blas:2009qj ; Anselmi:2008fk ; Barvinsky:2015kil ; Barvinsky:2021ubv ; Barvinsky:2023uir . The second—and in some sense the minimal one—is to keep the standard two–derivative structure (so that the canonical scaling is still relativistic, $z=1$) but to allow different fields, or different linear combinations of fields, to propagate on different effective “light cones” (or, in Euclidean language, with different positive–definite quadratic forms in spatial gradients and/or time derivatives). This second class of models is ubiquitous: it appears as the low–energy theory of coupled order parameters with different stiffnesses, as multi–field EFTs with distinct sound speeds Garriga:1999vw ; Cheung:2007st ; Weinberg:2008hq , and as toy models for “multi–metric” matter sectors in Lorentz–violating extensions of relativistic QFT Coleman:1998ti ; Jacobson:2000xp .
@@
Lorentz violation has been systematized from several complementary angles. In the EFT approach, the Standard–Model Extension provides a general local operator basis Kostelecky:1989jp ; Colladay:1996iz ; Colladay:1998fq ; Kostelecky:2003fs , while simplified “different limiting velocities” parametrizations were emphasized early on in Coleman:1998ti (see also classic LV electrodynamics examples such as Carroll:1989vb ; more about LV in such setups maybe also found in Rubtsov:2012kb ; Satunin:2017wmk ). Renormalization and operator mixing in LV gauge theories have been investigated in a number of sectors, including one–loop analyses in Lorentz–violating QED–type settings Kostelecky:2001jc . Reviews and continually updated experimental constraints can be found in Mattingly:2005re ; Jacobson:2005bg ; Liberati:2013xla ; Kostelecky:2008ts . Naturalness issues for small LV under radiative corrections have been highlighted in Myers:2003fd ; Collins:2004bp , with possible protection mechanisms in supersymmetric settings GrootNibbelink:2004za . On the many–body side, early RG arguments for the emergence of a common effective light cone go back to Chadha:1982ex (see also Volovik:2003fe ; Wen:2004ym ), and in Lorentz–violating gravity, controlled preferred–frame frameworks include Einstein–aether and Hořava–Lifshitz models Jacobson:2000xp ; Jacobson:2007ve ; Horava:2009uw ; Blas:2009qj ; Sotiriou:2009gy .
@@
Moreover, setups where scalar perturbations may propagate with reduced or field-dependent sound speeds arise in single- and multi-field inflation Cheung:2007st ; Armendariz-Picon:1999hyi ; Kobayashi:2010cm (and not only inflation Mironov:2018oec ; Ageeva:2022asq ), so that study of such theories with several scalar fields with different sound speeds was proposed in Refs.Ageeva:2022byg ; Ageev:2025fqv , where authors consider partial wave unitarity.
@@
The present work complements these by providing a compact, basis–covariant one–loop renormalization technology tailored to multi–metric two–derivative scalar theories with direction–dependent mixing. The renormalization in the presence of such spatial anisotropy in quantum field theory has been previously considered, e.g., in Refs. Arefeva:1994qr ; Periwal:1995wm .
@@
A basic structural question in this setting is: how do interactions renormalize when the quadratic kinetic operator is anisotropic and mixes fields in a direction–dependent way? Even for scalar theories, a fully systematic answer requires some care. First, the propagator is matrix–valued in field space, and kinetic mixing terms imply that the normal modes are not the original fields but momentum–dependent linear combinations. Second, in dimensional regularization and minimal subtraction Bollini:1972ui ; Hooft:1972fi ; Hooft:1973mf one must separate UV poles from possible IR subtleties of massless vacuum diagrams. Third, if one wants results that are robust under field redefinitions, it is desirable to express UV divergences not in a particular diagonal basis but directly in terms of invariant spectral data of the quadratic operator Kamefuchi:1961sb .
@@
One might naively think just to just introduce anisotropy in spatial derivatives and study such theory with kinetic term
@@
However to keep the theory RG complete (in the presence of parity-violating couplings) ha must consider more complicated version of such theory with addition of cross derivative term and keeping arbtitrary normalization, i.e. add to the theory terms of the form
$\frac{1}{2}\left(\partial_{0}\bm{\varphi}\right)^{\top}\mathbf{Z}_{t}\left(\partial_{0}\bm{\varphi}\right)+(\nabla\phi)^{\top}\mathbf{Y}(\nabla\chi).$
In this work we develop a one–loop renormalization of the theory for the simplest interacting model that contains all of these ingredients: two real scalar fields $\phi$ and $\chi$ in $D=4-\varepsilon$ Euclidean dimensions with the most general quadratic form built from two derivatives, including time–derivative mixing and cross–gradient mixing, together with the most general cubic and quartic interactions.
@@
We show that the one–loop UV poles of the theory can be written in a remarkably universal form. At $D=4-\epsilon$ all one–loop divergences are controlled by two “master integrals”: the tadpole (which carries one insertion of the mass operator in MS) and the bubble at zero external momentum. After performing the frequency integral, both reduce to angular averages over $\hat{\bm{n}}$ dressed by a single positive kernel
@@
    General one–loop MS renormalization for a two–scalar anisotropic QFT. We compute the complete set of one–loop UV divergences in dimensional regularization with minimal subtraction near $D=4-\epsilon$ for the theory (2)–(12). All counterterms and beta functions are expressed in compact tensor form in terms of two finite objects: the tadpole residue $\mathcal{T}_{ij}$ and the bubble residue $\mathcal{B}_{ij;kl}$, defined by $\mathcal{T}\equiv\epsilon\,\mathbb{T}$ and $\mathcal{B}\equiv\epsilon\,\mathbb{B}(0)$.
@@
    Locality of one–loop poles and non–running of anisotropy matrices in MS. A key structural simplification is that in $D=4-\epsilon$ the one–loop UV poles in two–point functions are momentum–independent (compare the standard Lorentz–invariant $\phi^{4}$ case, where wavefunction renormalization starts at two loops in MS; see e.g. ZinnJustin:2002ru ; Amit:2005cd ). As a consequence, in DR+MS at one loop
@@
    while the couplings $(\lambda_{ijkl},h_{ijk},M^{2}_{ij},j_{i})$ do run. In other words, at this order the anisotropy data act as external parameters that weight the RG flow of interactions but do not themselves evolve in MS. (Finite parts, other schemes, or higher loops can induce running of these matrices; our formalism is designed to extend to those settings.)
@@
    $\beta_{\lambda_{ijkl}}$ $=-\epsilon\,\lambda_{ijkl}+$ (4)
@@
    and obeys the universal inequality $J_{12}^{2}\leq J_{11}J_{22}$. Introducing the mismatch ratio $\rho\equiv J_{12}/\sqrt{J_{11}J_{22}}\in(0,1]$ makes many RG statements manifest.
@@
    Fixed points, existence bounds, and cubic couplings. In the parity–symmetric quartic sector ($\lambda_{4,5}=0$) we find coupled fixed points generalizing the standard two–scalar Wilson–Fisher point Wilson:1971dc ; Wilson:1971vs ; WilsonFisher:1971dc and the coupled–order–parameter multicritical analyses in $O(n_{1})\oplus O(n_{2})$–type theories Aharony:1973zz ; Pelissetto:2000ek ; Calabrese:2002bm . Their very existence depends on anisotropy through $\rho$: the IR–stable coupled point exists only for sufficiently small mismatch, $\rho\geq\sqrt{3}/2$ (equivalently $4J_{12}^{2}\geq 3J_{11}J_{22}$), and disappears when the mismatch is too large. In the fully coupled quartic sector ($\lambda_{4,5}\neq 0$) the one–loop equations admit a fixed line (a marginal direction at this order), which we analyze and show to be of saddle type. Cubic couplings are relevant already at tree level in $4-\epsilon$ dimensions, and at one loop the cubic flow remains linear in $h$ (see e.g. Amit:2005cd ; ZinnJustin:2002ru ). In particular, within the perturbative $4-\epsilon$ expansion there is no interacting fixed point with $h^{\star}\neq 0$; cubic perturbations destabilize the quartic critical loci unless they are tuned to zero (or forbidden by a symmetry).
@@ 2 Setup: theory and its decomposition
In this paper we consider the theory of two real interacting scalar fields $\phi$ and $\chi$ in $D\equiv d+1$ space-time and stick to Euclidean signature of the metric. The Lagrangian under consideration is given by
@@
where [^1] Equivalently, the quadratic derivative terms can be written in index notation as $\mathcal{L}_{kin}=\frac{1}{2}\,Z_{t}^{ij}\,\partial_{0}\varphi_{i}\,\partial_{0}\varphi_{j}+\frac{1}{2}\,C_{\phi}^{ab}\,\partial_{a}\phi\,\partial_{b}\phi+\frac{1}{2}\,C_{\chi}^{ab}\,\partial_{a}\chi\,\partial_{b}\chi+Y^{ab}\,\partial_{a}\phi\,\partial_{b}\chi$ $\bm{\varphi}\equiv(\phi,\chi)^{\top}$ and the interactions have the form
@@
The counterterms for renormalization in this theory has the form
$\mathcal{L}_{\mathrm{ct}}=\frac{1}{2}(\partial_{0}\bm{\varphi})^{\top}\delta\mathbf{Z}_{t}(\partial_{0}\bm{\varphi})+\frac{1}{2}(\nabla\phi)^{\top}\delta\mathbf{C}_{\phi}(\nabla\phi)+$ (14)
$+\frac{1}{2}(\nabla\chi)^{\top}\delta\mathbf{C}_{\chi}(\nabla\chi)+(\nabla\phi)^{\top}\delta\mathbf{Y}(\nabla\chi)+\frac{1}{2}\,\bm{\varphi}^{\top}\delta\mathbf{M}^{2}\,\bm{\varphi}+$ (15)
$+\mu^{\frac{D+2}{2}}\,\delta j_{i}\,\varphi_{i}+\frac{\mu^{\frac{6-D}{2}}}{3!}\,\delta h_{ijk}\,\varphi_{i}\varphi_{j}\varphi_{k}+\frac{\mu^{4-D}}{4!}\,\delta\lambda_{ijkl}\,\varphi_{i}\varphi_{j}\varphi_{k}\varphi_{l}$ (16)
@@
  • $\mathbf{Z}_{t}$ is a $2\times 2$ positive definite which is necessary for well‑posed Euclidean path integral and stability. This matrix is defined as
@@
  • $C_{\phi},C_{\chi}$ are “sound‑speed” [^2] We tacitly assume that they are sound-speed matrices in Lorentzian signature or control; in the Euclidean signature could be called stiffness matrices. real symmetric positive‑definite $d\times d$ matrices acting on spatial indices. For example, if one considers $d=2$ case, then $C_{\phi}$ is
    $C_{\phi}=\begin{pmatrix}C^{(\phi)}_{xx}\;\;\;C^{(\phi)}_{xy}\\ C^{(\phi)}_{xy}\;\;\;C^{(\phi)}_{yy}\end{pmatrix}$. (18)
  • $\mathbf{Y}$ is a real symmetric $d\times d$ matrix that mixes spatial gradients i.e. $Y^{ab}\,\partial_{a}\phi\,\partial_{b}\chi$.
@@
We keep the most general quadratic form, including unusual terms like $\mathbf{Z}_{t}$ and $\mathbf{Y}$, because it is closed under field redefinitions and expected to be generated beyond one loop or in other dimensional expansions. In the present one-loop DR+MS analysis near $D=4-\varepsilon$, the UV $1/\varepsilon$ poles in two-point functions are momentum independent, hence they renormalize only $\mathbf{M}^{2}$ and not $\mathbf{Z}_{t}$, $\mathbf{C}_{\phi,\chi}$, or $\mathbf{Y}$. The form of counter terms matrices $\delta\mathbf{Z}_{t}$, $\delta C_{\phi,\chi}$, $\delta Y$, and $\delta M_{ij}^{2}$ is defined by the term which this counterterm renormalize.
Next we turn to the interaction part of the renormalized Lagrangian (13). Here we have the totally symmetric tensors $h_{ijk}=h_{(ijk)}$ and $\lambda_{ijkl}=\lambda_{(ijkl)}$, where $i,j,k,l\in\{1,2\}$. If we write (13) explicitly, we arrive to expression (12) where tensor notations of (13) is related to (12) as
@@
with $\mathbf{Z}_{t}^{1/2}$ being a square root of the matrix $\mathbf{Z}_{t}$, which is well-defined since $\mathbf{Z}_{t}$ is symmetric and positive-define, as it was mentioned above. Here $\mathbf{1}$ denotes the $2\times 2$ identity in field space. We take $\mathbf{Z}_{t}^{1/2}$ to be the principal symmetric square root, so that $\mathbf{Z}_{t}^{1/2}=\mathbf{Z}_{t}^{1/2\,\top}$.
@@
$\tilde{\bm{\Pi}}_{a}(\mathbf{p})\equiv\mathbf{u}_{a}(\mathbf{p})\mathbf{u}_{a}(\mathbf{p})^{T},\quad\sum_{a=1}^{2}\tilde{\bm{\Pi}}_{a}=\mathbf{I},\quad\tilde{\bm{\Pi}}_{a}\tilde{\bm{\Pi}}_{b}=\delta_{ab}\tilde{\bm{\Pi}}_{a}.$ (32)
@@
$\bm{\Pi}_{a}(\hat{\mathbf{n}})\equiv\mathbf{v}_{a}(\hat{\mathbf{n}})\mathbf{v}_{a}(\hat{\mathbf{n}})^{T},\quad\sum_{a=1}^{2}\bm{\Pi}_{a}=\mathbf{I},\quad\bm{\Pi}_{a}\bm{\Pi}_{b}=\delta_{ab}\bm{\Pi}_{a},$ (39)

### v2

# INR-TH-2025-024

@@ Abstract
We develop a basis–covariant one–loop renormalization framework for two interacting real scalars in $D=4-\epsilon$ with the most general two–derivative Lorentz–violating quadratic form, allowing anisotropic spatial gradients and direction–dependent kinetic mixing, together with general cubic and quartic interactions forming RG complete set of operators at one-loop. In dimensional regularization with minimal subtraction we compute the full set of one–loop UV divergences and obtain closed beta functions for quartic and cubic couplings, masses.
The pole coefficients admit a universal spectral representation as angular averages over the direction–dependent eigenvalues and projectors of the UV kinetic matrix; all anisotropy dependence enters through a single universal kernel admitting two–particle phase–space interpretation. We classify fixed points and fixed manifolds and show, in particular, that anisotropy restricts the existence of the coupled Wilson–Fisher–type fixed point. When the cross–gradients are turned on the coefficients in beta functions are governed by six phase–space weights admitting interpretation in terms of encoding “populations” and “coherences” of the UV normal modes.
@@ 1 Introduction
Lorentz invariance is one of the organizing principles of high–energy physics: it underlies the locality and analyticity properties of relativistic quantum field theory (QFT), it severely constrains the spectrum of operators, and provides the kinematic backbone for essentially all perturbative calculations. At the same time, there are many reasons—both conceptual and practical—to study controlled departures from exact Lorentz symmetry. On the one hand, Lorentz invariance may be an emergent symmetry of an interacting many–body system at long distances: in condensed–matter realizations, different fields typically propagate with different “speeds” and the ultraviolet (UV) theory is anisotropic, while approximate relativistic invariance can appear only in the deep infrared (IR) region Chadha:1982ex ; Volovik:2003fe ; Wen:2004ym . On the other hand, in particle physics and quantum gravity it is natural to ask how robust the Lorentz–invariant fixed points are under Lorentz–violating (LV) perturbations, and whether small violations can remain technically natural under renormalization Myers:2003fd ; Collins:2004bp ; GrootNibbelink:2004za (as explored, for instance, in effective field theory’s (EFT) frameworks such as the Standard–Model Extension and its gravitational generalizations Kostelecky:1989jp ; Colladay:1996iz ; Colladay:1998fq ; Kostelecky:2003fs ; Kostelecky:2008ts and in the broader phenomenological literature Mattingly:2005re ; Jacobson:2005bg ; Liberati:2013xla ).
@@
There are (at least) two complementary ways in which Lorentz invariance can be violated in a continuum field theory. The first is to change the scaling structure by introducing higher spatial derivatives (Lifshitz–type theories with dynamical exponent $z\neq 1$ [^1] See for example Kachru:2008yh and references therein., including the Hořava–Lifshitz construction in gravity) Lifshitz:1941 ; Horava:2009uw ; Sotiriou:2009gy ; Blas:2009qj ; Anselmi:2008fk ; Barvinsky:2015kil ; Barvinsky:2021ubv ; Barvinsky:2023uir . The second—and in some sense the minimal one—is to keep the standard two–derivative structure (so that the canonical scaling is still relativistic, $z=1$) but to allow different fields, or different linear combinations of fields, to propagate on different effective “light cones” (or, in Euclidean language, with different positive–definite quadratic forms in spatial gradients and/or time derivatives). This second class of models is ubiquitous: it appears as the low–energy theory of coupled order parameters with different stiffnesses, as multi–field EFTs with distinct sound speeds Garriga:1999vw ; Cheung:2007st ; Weinberg:2008hq , and as toy models for “multi–metric” matter sectors in Lorentz–violating extensions of relativistic QFT Coleman:1998ti ; Jacobson:2000xp .
@@
Lorentz violation has been systematized from several complementary angles. In the EFT approach, the Standard–Model Extension provides a general local operator basis Kostelecky:1989jp ; Colladay:1996iz ; Colladay:1998fq ; Kostelecky:2003fs , while simplified “different limiting velocities” parametrizations were emphasized in Coleman:1998ti ; Iengo:2009ix ; Anber:2011xf . See also classic LV electrodynamics examples such as Carroll:1989vb ; more about LV in such setups maybe also found in Rubtsov:2012kb ; Satunin:2017wmk . Renormalization and operator mixing in LV gauge theories have been investigated in a number of sectors, including one–loop analysis in Lorentz–violating scalar-spinor interactions Altschul:2006jj ; Ferrero:2011yu ; Altschul:2012xu ; Altschul:2014gqa , quantum electrodynamic (QED) – type settings Kostelecky:2001jc ; Altschul:2022isc and quantum chronodynamics (QCD) – in Altschul:2023vnf . Reviews and continually updated experimental constraints can be found in Mattingly:2005re ; Jacobson:2005bg ; Liberati:2013xla ; Kostelecky:2008ts . Naturalness issues for small LV under radiative corrections have been highlighted in Myers:2003fd ; Collins:2004bp , with possible protection mechanisms in supersymmetric settings GrootNibbelink:2004za . On the many–body side, early renormalization group (RG) arguments for the emergence of a common effective light cone go back to Chadha:1982ex (see also Volovik:2003fe ; Wen:2004ym ), and in Lorentz–violating gravity, controlled preferred–frame frameworks include Einstein–aether and Hořava–Lifshitz models Jacobson:2000xp ; Jacobson:2007ve ; Horava:2009uw ; Blas:2009qj ; Sotiriou:2009gy . In Affleck:2021vzo authors consider the renormalization of velocities in non-Lorentz invariant sigma model, showing that the velocity differences flow to zero at low enough energies and it is possible to restore the Lorentz invariance.
Related setups – where scalar perturbations propagate with reduced or field‑
@@
dependent sound speeds – arise in single‑ and multi‑field inflation Cheung:2007st ; Armendariz-Picon:1999hyi ; Kobayashi:2010cm and in other contexts Mironov:2018oec ; Ageeva:2022asq . The study of multi‑scalar theories with different sound speeds was proposed in Refs. Ageeva:2022byg ; Ageev:2025fqv , where the authors analyze partial‑wave unitarity.
@@
The present work complements these studies by providing a compact, basis–covariant one–loop renormalization machinery tailored to multi–metric two–derivative scalar theories with direction–dependent mixing. Related renormalization analyses for spatially anisotropic QFTs include Refs. Arefeva:1994qr ; Periwal:1995wm . Another similar setup is $O(N)$ self-interacting scalar field theory with Lorentz violation from Carvalho:2013wsa , where a mass renormalization up to two-loop level has been proceeded.
@@
A basic structural question in this setting is: how do interactions renormalize when the quadratic kinetic operator is anisotropic and mixes fields in a direction–dependent way? Even for scalar theories, a fully systematic answer requires some care. First, the propagator is matrix–valued in field space, and kinetic mixing terms imply that the normal modes are not the original fields but momentum–dependent linear combinations. For example, in Bijnens:2018rqw there is a similar to ours setup with both kinetic and mass mixing: it was investigated whether the kinetic term can be kept diagonal under renormalization group evolution, even though the fields mix. Second, in dimensional regularization and minimal subtraction Bollini:1972ui ; Hooft:1972fi ; Hooft:1973mf one must separate UV poles from possible IR subtleties of massless vacuum diagrams. Third, if one wants results that are robust under field redefinitions, it is desirable to express UV divergences not in a particular diagonal basis but directly in terms of invariant spectral data of the quadratic operator Kamefuchi:1961sb .
@@
One might naively think just to introduce anisotropy in spatial derivatives and study such theory with kinetic term as follows
@@
However to keep the theory RG complete (in the presence of parity-violating couplings) one must consider more complicated version of such theory with addition of cross derivative term and keeping arbitrary normalization, i.e. add to the theory terms of the form
$\frac{1}{2}\left(\partial_{0}\bm{\varphi}\right)^{\top}\mathbf{Z}_{t}\left(\partial_{0}\bm{\varphi}\right)+(\nabla\phi)^{\top}\mathbf{Y}(\nabla\chi).$ (1)
The latter term with gradient mixing may be not optional and arises in different physical systems, for example in superfluids solutions Andreev . In this work we develop a one–loop renormalization of the theory for the simplest interacting model that contains all of these ingredients: two real scalar fields $\phi$ and $\chi$ in $D=4-\epsilon$ Euclidean dimensions with the most general quadratic form built from two derivatives, including time–derivative mixing and cross–gradient mixing, together with the most general cubic and quartic interactions.
@@
We show that the one–loop UV poles of the theory can be written in a remarkably universal form. At $D=4-\epsilon$ all one–loop divergences are controlled by two “master integrals”: the tadpole (which carries one insertion of the mass operator in minimal subtraction (MS) renormalization scheme) and the bubble at zero external momentum. After performing the frequency integral, both reduce to angular averages over $\hat{\bm{n}}$ dressed by a single positive kernel
@@
    General one–loop MS renormalization for a two–scalar anisotropic QFT. We compute the complete set of one–loop UV divergences in dimensional regularization with minimal subtraction near $D=4-\epsilon$ for the theory (2)–(12). All counterterms and beta functions are expressed in compact tensor form in terms of two finite objects: the tadpole residue $\mathcal{T}_{ij}$ and the bubble residue $\mathcal{B}_{ij;kl}$.
@@
    Locality of one–loop poles and non–running of anisotropy matrices in MS. A key structural simplification is that in $D=4-\epsilon$ the one–loop UV poles in two–point functions are momentum–independent (compare the standard Lorentz–invariant $\phi^{4}$ case, where wavefunction renormalization starts at two loops in MS; see e.g. ZinnJustin:2002ru ; Amit:2005cd ). As a consequence, in dimensional regularization (DR)+MS at one loop
@@
    while the couplings $(\lambda_{ijkl},h_{ijk},M^{2}_{ij},j_{i})$ do run. Notice, that while there are no divergent parts to get non-vanishing beta functions finite pieces may be generated at one-loop level and this is the reason to include them obtaining RG closed set of operators. In other words, at this order the anisotropy data act as external parameters that weight the RG flow of interactions but do not themselves evolve in MS. Finite parts, other schemes, or higher loops can induce running of these matrices; our formalism is designed to extend to those settings.
@@
    $\beta_{\lambda_{ijkl}}$ $=-\epsilon\,\lambda_{ijkl}+$
@@
    and obeys the universal inequality $J_{12}^{2}\leq J_{11}J_{22}$. Introducing the mismatch ratio $\rho\equiv J_{12}/\sqrt{J_{11}J_{22}}\in(0,1]$ makes many RG statements manifest. In the RG closed sector corresponding to parity-symmetric quartic couplings case beta functions have simple form
    $\beta_{\lambda_{1}}$ $=-\epsilon\lambda_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}^{2}+J_{22}\lambda_{3}^{2}\right],$
    $\beta_{\lambda_{2}}$ $=-\epsilon\lambda_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}^{2}+J_{11}\lambda_{3}^{2}\right],$
    $\beta_{\lambda_{3}}$ $=-\epsilon\lambda_{3}+\frac{1}{16\pi^{2}}\left[J_{11}\lambda_{1}\lambda_{3}+4J_{12}\lambda_{3}^{2}+J_{22}\lambda_{2}\lambda_{3}\right],$
    structurally similar to the isotropic case.
@@
    Fixed points, existence bounds, and cubic couplings. In the parity–symmetric quartic sector ($\lambda_{4,5}=0$) we find coupled fixed points generalizing the standard two–scalar Wilson–Fisher point Wilson:1971dc ; Wilson:1971vs ; WilsonFisher:1971dc and the coupled–order–parameter multicritical analyses in $O(n_{1})\oplus O(n_{2})$–type theories Aharony:1973zz ; Pelissetto:2000ek ; Calabrese:2002bm . Their very existence depends on anisotropy through $\rho$: the IR–stable coupled point exists only for sufficiently small mismatch, $\rho\geq\sqrt{3}/2$ (equivalently $4J_{12}^{2}\geq 3J_{11}J_{22}$), and disappears when the mismatch is too large. In the fully coupled quartic sector ($\lambda_{4,5}\neq 0$) the one–loop equations admit a fixed line (a marginal direction at this order), which we analyze and show to be of saddle type. Cubic couplings are relevant already at tree level in $4-\epsilon$ dimensions, and at one loop the cubic flow remains linear in $h$ (see e.g. Amit:2005cd ; ZinnJustin:2002ru ). In particular, within the perturbative $4-\epsilon$ expansion there is no interacting fixed point with $h^{*}\neq 0$; cubic perturbations destabilize the quartic critical loci unless they are tuned to zero (or forbidden by a symmetry).
@@ 2 Setup: theory and its decomposition
In this paper we consider the theory of two real interacting scalar fields $\phi$ and $\chi$ in $D\equiv d+1$ Euclidean spacetime dimensions. The Lagrangian under consideration is given by
@@
where [^2] Equivalently, the quadratic derivative terms could be written in index notation as $\mathcal{L}_{kin}=\frac{1}{2}\,Z_{t}^{ij}\,\partial_{0}\varphi_{i}\,\partial_{0}\varphi_{j}+\frac{1}{2}\,C_{\phi}^{ab}\,\partial_{a}\phi\,\partial_{b}\phi+\frac{1}{2}\,C_{\chi}^{ab}\,\partial_{a}\chi\,\partial_{b}\chi+Y^{ab}\,\partial_{a}\phi\,\partial_{b}\chi$. $\bm{\varphi}\equiv(\phi,\chi)^{\top}$ and the interactions have the form
@@
The counterterms for renormalization in this theory have the form
$\mathcal{L}_{\mathrm{ct}}=\frac{1}{2}(\partial_{0}\bm{\varphi})^{\top}\delta\mathbf{Z}_{t}(\partial_{0}\bm{\varphi})+\frac{1}{2}(\nabla\phi)^{\top}\delta\mathbf{C}_{\phi}(\nabla\phi)+$
$+\frac{1}{2}(\nabla\chi)^{\top}\delta\mathbf{C}_{\chi}(\nabla\chi)+(\nabla\phi)^{\top}\delta\mathbf{Y}(\nabla\chi)+\frac{1}{2}\,\bm{\varphi}^{\top}\delta\mathbf{M}^{2}\,\bm{\varphi}+$
$+\mu^{\frac{D+2}{2}}\,\delta j_{i}\,\varphi_{i}+\frac{\mu^{\frac{6-D}{2}}}{3!}\,\delta h_{ijk}\,\varphi_{i}\varphi_{j}\varphi_{k}+\frac{\mu^{4-D}}{4!}\,\delta\lambda_{ijkl}\,\varphi_{i}\varphi_{j}\varphi_{k}\varphi_{l}.$ (14)
@@
  • $\mathbf{Z}_{t}$ is a $2\times 2$ positive definite matrix in field space. Positivity is required for a well‑defined Euclidean path integral and for stability. This matrix is defined as
@@
  • $\mathbf{C}_{\phi},\mathbf{C}_{\chi}$ are “sound‑speed” [^3] We tacitly assume that they are sound-speed matrices in Lorentzian signature; in the Euclidean signature they could be called stiffness matrices. real symmetric positive‑definite $d\times d$ matrices acting on spatial indices. For example, if one considers $d=2$ case, then $\mathbf{C}_{\phi}$ is
    $\mathbf{C}_{\phi}=\begin{pmatrix}C^{(\phi)}_{xx}\;\;\;C^{(\phi)}_{xy}\\ C^{(\phi)}_{xy}\;\;\;C^{(\phi)}_{yy}\end{pmatrix}$. (16)
  • $\mathbf{Y}$ is a real symmetric $d\times d$ matrix that mixes spatial gradients, i.e. $Y^{ab}\,\partial_{a}\phi\,\partial_{b}\chi$.
@@
We keep the most general quadratic form, including unusual terms like $\mathbf{Z}_{t}$ and $\mathbf{Y}$, because it provides RG closed basis at least at one loop level.
[…]
In the present one-loop DR+MS analysis near $D=4-\varepsilon$, the UV $1/\varepsilon$ poles in two-point functions are momentum independent, hence they renormalize only $\mathbf{M}^{2}$ and not $\mathbf{Z}_{t}$, $\mathbf{C}_{\phi,\chi}$, or $\mathbf{Y}$. The form of counterterms matrices $\delta\mathbf{Z}_{t}$, $\delta\mathbf{C}_{\phi,\chi}$, $\delta\mathbf{Y}$, and $\delta M_{ij}^{2}$ is defined by the term which this counterterm renormalize.
Next we turn to the interaction part of the renormalized Lagrangian (13). Here we have the totally symmetric tensors $h_{ijk}=h_{(ijk)}$ and $\lambda_{ijkl}=\lambda_{(ijkl)}$, where $i,j,k,l\in\{1,2\}$. If we write (13) explicitly, we arrive at expression (12) where tensor notations of (13) is related to (12) as
@@
with $\mathbf{Z}_{t}^{1/2}$ being a square root of the matrix $\mathbf{Z}_{t}$, which is well-defined since $\mathbf{Z}_{t}$ is symmetric and positive-definite, as it was mentioned above. Here $\mathbf{1}$ denotes the $2\times 2$ identity in field space. We take $\mathbf{Z}_{t}^{1/2}$ to be the principal symmetric square root, so that $\mathbf{Z}_{t}^{1/2}=\mathbf{Z}_{t}^{1/2\,\top}$.
@@
$\tilde{\bm{\Pi}}_{a}(\mathbf{p})\equiv\mathbf{u}_{a}(\mathbf{p})\mathbf{u}_{a}(\mathbf{p})^{T},\quad\sum_{a=1}^{2}\tilde{\bm{\Pi}}_{a}=\mathbf{1},\quad\tilde{\bm{\Pi}}_{a}\tilde{\bm{\Pi}}_{b}=\delta_{ab}\tilde{\bm{\Pi}}_{a}.$ (30)
@@
$\bm{\Pi}_{a}(\hat{\mathbf{n}})\equiv\mathbf{v}_{a}(\hat{\mathbf{n}})\mathbf{v}_{a}(\hat{\mathbf{n}})^{T},\quad\sum_{a=1}^{2}\bm{\Pi}_{a}=\mathbf{1},\quad\bm{\Pi}_{a}\bm{\Pi}_{b}=\delta_{ab}\bm{\Pi}_{a},$ (37)